Under root x+y=7
X+under rooty=11,find the value of x and y

Answer by rothauserc(4718)

square root(x+y) = 7
x + square root(y) = 11
***********************************************
square both sides of = in first equation
x+y = 49
y = 49 - x
***********************************************
second equation subtract x from both sides of =
square root(y) = 11 - x
square both sides of =
y = 121 - 22x + x^2
***********************************************
we have two equations in two unknowns
y = 49 -x
y = 121 - 22x + x^2
substitute for y in second equation
49 - x = 121 - 22x + x^2
x^2 - 21x + 72 = 0
use quadratic formula to solve for x
x = (21 + 12.369316877) / 2 = 16.684658439
x = (21 - 12.369316877) / 2 = 4.315341562
our solution is
x = 16.684658439 or x = 4.315341562